OpenStudy (anonymous):

The angle of elevation of a ladder leaning against a wall is 45°. If the distance from the base of the ladder to the wall is 12 feet, how long is the ladder?

jhonyy9 (jhonyy9):

so can you draw it ?

OpenStudy (anonymous):

doesnt that just create a 45 90 45 degree triangle with the ladder as the hypotnus?

OpenStudy (kropot72):

cos 45 degrees = 12/h where h is the length of the ladder h = 12/cos 45 = 16.97 feet

OpenStudy (anonymous):

you have the angle against the wall (45) and the length of the hypotenuse (25ft) and you want to find the adajcent side. so for this we need to use 'cos' (CAH = cos(angle) =adjacent/hypotenuse) therefore: cos(45)=?/25 rearrange to get 25cos(45)=? type into calculator to get 16.73 fee

OpenStudy (anonymous):

\[\cos(45)=?/25 rearrange \to get 25\cos(45)=? type into calculator \to get 16.73 fee \]

OpenStudy (anonymous):



OpenStudy (anonymous):

@iceywheniplease7 see ^^

OpenStudy (anonymous):

that's a nice triangle

OpenStudy (anonymous):

ohh now i got it! than ks! :)